# Training-Conditional Coverage Bounds under Covariate Shift

### v1

@@ C.2 Proof of Theorem 3
$P_{e}({\bf D}_{n})=\operatorname{\mathbb{P}}(Y_{n+1}\notin\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})|{\bf D}_{n})$ $\leq\operatorname{\mathbb{P}}\left(1-F_{Q}^{(n-1)}\left(\left|Y_{n+1}-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+2\epsilon L_{Q,n-1}\right)$
@@
$\operatorname{\mathbb{P}}({\bf D}_{n}\notin\mathcal{A})\leq 2p\exp\left(-\frac{2\kappa_{1}^{2}}{n}\left(\frac{\epsilon}{\kappa_{2}c_{n-1}}-\frac{1}{\kappa_{1}}\right)^{2}\right)$
@@
$\operatorname{\mathbb{P}}\left({\bf D}_{n}\notin\mathcal{B}\left(\delta+3\,C\sqrt{\frac{B}{n}}\right)\right)\leq 4e^{-n\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta+\epsilon+3\,C\sqrt{\frac{B}{n}}\right)\leq 4$ $e^{-n\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log\frac{4}{\delta}}+3C\right)\sqrt{\frac{B}{n}}+2\,L_{Q,n-1}\,\kappa_{2}\,c_{n-1}$ $\left(\frac{1}{\kappa_{1}}+\sqrt{\frac{n}{2\kappa_{1}^{2}}\log\frac{2p}{\epsilon}}\right)\Bigg{)}$
@@ D.1 Proof of Theorem 4
$P_{e}({\bf D}_{n})$ $=\operatorname{\mathbb{P}}(Y_{n+1}\notin\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})|{\bf D}_{n})$
$\leq\operatorname{\mathbb{P}}\left(1-F^{(n)}\left(\left|Y_{n+1}-\mu_{\overline{\beta}_{n}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+(2\epsilon+c_{n+1})L_{n}\right)$
@@ D.2 Proof of Theorem 5
$P_{e}({\bf D}_{n})$ $=\operatorname{\mathbb{P}}(Y_{n+1}\notin\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})|{\bf D}_{n})$
$\leq\operatorname{\mathbb{P}}\left(1-F_{Q}^{(n)}\left(\left|Y_{n+1}-\mu_{\overline{\beta}_{n}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+(2\epsilon+c_{n+1})L_{Q,n}\right)$
@@
$\operatorname{\mathbb{P}}\left({\bf D}_{n}\notin\mathcal{B}\left(\delta+3\,C\sqrt{\frac{B}{n}}\right)\right)\leq 4e^{-n\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta+3\,C\sqrt{\frac{B}{n}}+\epsilon\right)\leq 4e^{-n\delta^{2}/(2B^{2})}+2p\exp\left(-\left(\frac{\kappa_{1}(\epsilon/L_{Q,n}-c_{n+1})}{\sqrt{2n}\kappa_{2}c_{n}}\right)^{2}\right),$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+3C\right)\sqrt{\frac{B}{n}}+L_{Q,n}\left(c_{n+1}+\sqrt{2n\log\frac{2p}{\epsilon}}\,\frac{\kappa_{2}\,c_{n}}{\kappa_{1}}\right)\right)\leq\epsilon+\delta.$

### v2

@@ C.2 Proof of Theorem 4
$P_{e}({\bf D}_{n})=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}^{\text{J+}}(X_{\text{test}})|{\bf D}_{n})$ $\leq\operatorname{\mathbb{P}}\left(1-F_{Q}^{(n-1)}\left(\left|Y_{\text{test}}-\mu_{\overline{\beta}_{-1}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+2\epsilon L_{Q,n-1}\right)$
@@
$\operatorname{\mathbb{P}}\left({\bf D}_{n}\notin\mathcal{B}\left(\delta+2C\sqrt{\frac{B}{n}}\right)\right)\leq 4e^{-n\delta^{2}/(2B^{2})}.$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta+\epsilon+2C\sqrt{\frac{B}{n}}\right)\leq 4$ $e^{-n\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\Bigg{(}P_{e}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log\frac{4}{\delta}}+2C\right)\sqrt{\frac{B}{n}}+2\,L_{Q,n-1}\,\kappa_{2}\,c_{n-1}$ $\left(\frac{1}{\kappa_{1}}+\sqrt{\frac{n}{2\kappa_{1}^{2}}\log\frac{2p}{\epsilon}}\right)\Bigg{)}$
@@ D.1 Proof of Theorem 5
$P_{e}({\bf D}_{n})$ $=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})|{\bf D}_{n})$
$\leq\operatorname{\mathbb{P}}\left(1-F^{(n)}\left(\left|Y_{\text{test}}-\mu_{\overline{\beta}_{n}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+(2\epsilon+c_{n+1})L_{n}\right)$
@@ D.2 Proof of Theorem 6
$P_{e}({\bf D}_{n})$ $=\operatorname{\mathbb{P}}(Y_{\text{test}}\notin\hat{\mathcal{C}}_{\alpha}^{\text{full}}(X_{\text{test}})|{\bf D}_{n})$
$\leq\operatorname{\mathbb{P}}\left(1-F_{Q}^{(n)}\left(\left|Y_{\text{test}}-\mu_{\overline{\beta}_{n}}(X_{\text{test}})\right|\right)\leq\alpha+\delta+(2\epsilon+c_{n+1})L_{Q,n}\right)$
@@
$\operatorname{\mathbb{P}}\left({\bf D}_{n}\notin\mathcal{B}\left(\delta+2C\sqrt{\frac{B}{n}}\right)\right)\leq 4e^{-n\delta^{2}/(2B^{2})}$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\delta+2C\sqrt{\frac{B}{n}}+\epsilon\right)\leq 4e^{-n\delta^{2}/(2B^{2})}+2p\exp\left(-\left(\frac{\kappa_{1}(\epsilon/L_{Q,n}-c_{n+1})}{\sqrt{2n}\kappa_{2}c_{n}}\right)^{2}\right),$
@@
$\operatorname{\mathbb{P}}\left(P_{e}({\bf D}_{n})>\alpha+\left(\sqrt{2B\log{4}/{\delta}}+2C\right)\sqrt{\frac{B}{n}}+L_{Q,n}\left(c_{n+1}+\sqrt{2n\log\frac{2p}{\epsilon}}\,\frac{\kappa_{2}\,c_{n}}{\kappa_{1}}\right)\right)\leq\epsilon+\delta.$